OpenStudy (anonymous):

Determine this quotient. (-2.8) ÷ 4

OpenStudy (alexandervonhumboldt2):

\[-\frac{ 28 }{ 10 }/4=-\frac{ 14 }{ 5 }*\frac{ 1 }{ 4 }=\frac{ -14*1 }{ 5*4 }=-\frac{ 14 }{ 20 }=-\frac{ \cancel 2*7 }{ \cancel 2*10 }=-7/10\]